The John Hancock Center in Chicago has a rectangular base. The length of the base measures 65 feet less than twice the width. The perimeter of this base is 860 feet. What are the dimensions of the base?
Answer by mananth(16946)

Perimeter = 860
Width =x ft
Length 2x-65
Perimeter = 2((2x-65)+x)
Perimeter = 2( 2x-65+x)
Perimeter = 4x-130+2x
860 =6x-130
860 + 130 = 6 x
990 = 6 x
x= 165 ft Width
Length = 265 ft